### v1

@@ 1 Introduction
Reinforcement learning (RL) is a data-driven approach to the problem of goal-directed sequential decision making. The RL framework involves an ‘agent’, which is its decision-making component, that learns a mapping from states to actions, called the ‘policy’, via a series of trial-and-error interactions with the environment. A reward signal, which is a scalar measure of performance of the environment in relation to the pre-defined goal, is fed back to the agent and is used in the optimization. The state $S_{t}$, action $A_{t}$ and reward $R_{t}$ form the three fundamental channels of communication between the agent and the environment. Each state is assigned a value $V(s)$ which is the expected total reward that can be obtained by that state in the long run, i.e., $V(s)=\operatorname*{\mathbb{E}}\left[\Sigma_{k=0}^{\infty}\gamma^{k}R_{t+k+1}|s\right]$ [47]. The goal of the RL problem, then, is to maximise expected cumulative rewards or ‘returns’. In Deep RL (DRL), function approximators such as neural networks are used to parameterise either the policy, the value function, or both, which are updated during training via interaction with the environment in order to maximise the returns. Various solution methods have been developed to perform this optimisation, namely: (i) value-based methods such as Q-learning method [58, 28], where the value function is approximated and actions are chosen based on the approximated value function, (ii) policy-based methods [50], such as the policy gradient method, where the optimal policy is learned directly, and (iii) actor-critic methods such as the proximal policy optimisation (PPO) method [53], which leverages the advantages of both methods. During training, interaction between the agent and the environment occurs on an episode basis, where an ‘episode’ is a single trajectory of the environment for a given fixed duration. Training is performed through trial-and-error that involves a stochastic exploration of the environment’s states, i.e., selecting actions other than those dictated by the latest updated policy, which can improve the estimation of the state values. After training, the agent is evaluated on the environment in a deterministic manner according to the optimal policy, without exploration.
@@
As a consequence, DRL for AFC has been the focus of much work over the last few years. One of the first uses of DRL in AFC was for the stabilisation of the vortex street in the flow past a two-dimensional (2D) cylinder at a low Reynolds number ($Re$) of 100 [38, 37]. Here, the Reynolds number is given as $Re=\tilde{U}\tilde{H}/\tilde{\nu}$, where the symbols $\tilde{U}$, $\tilde{H}$ and $\tilde{\nu}$ are the mean flow velocity, characteristic length scale of the system and kinematic viscosity respectively, and tilde $\tilde{\cdot}$ represents dimensional quantities. With two synthetic jets as the actuators, the drag was reduced by $8\%$ and the learned control strategy was found to reduce pressure drop across the cylinder by increasing the area of the separated region. Following this study, DRL was applied to different configurations of the 2D cylinder. In Tang et al. [56], four synthetic jets were used as DRL controllers for a wider $Re$ range, $Re=100-400$, and a maximum drag reduction of $\approx 38\%$ was achieved. Drag reduction was observed for unseen values of $Re$ as well, showing the robustness of the DRL control process. The work by Ren et al. [39] extends these studies to weakly turbulent conditions at $Re=1000$, showing a reduction in drag by $30\%$ along with a reduction of turbulent fluctuations. Varela et al. [63] shows that a transition in the actuation strategy is obtained when the Reynolds number is further increased and turbulent stresses play a dominating role. Fan et al. [16] and Xu et al. [68] use a pair of smaller downstream rotating cylinders as actuators with a combined objective of drag reduction over the main cylinder and also to reduction of power loss due to cylinder rotation. More recently, an application to three-dimensional (3D) cylinders was studied in the $Re$ range of $100-400$ [45], where again, DRL successfully showed a reduction of drag of $\approx 10\%$ for $Re=400$. A number of studies have also considered questions related to the optimization of the full DRL control setup, including probe placement [34, 33, 69] and the use of time history from the state probes [66]. Moreover, applications related to vortex induced vibration [41, 12, 44], or hydrodynamic stealth [42], have also attracted interest. Aside from the cylinder, other applications include the control of chaotic flows such as the flow arising from the Kuramoto–Sivashinsky (KS) equation [9, 67], drag reduction in 3D channel flows [19, 51], suppression of the instability of a one-dimensional (1D) falling liquid film [8], suppression of 2D Rayleigh–Bénard convection (RBC) [3, 62] and other practical problems [41, 42, 27]. We refer the interested reader to some comprehensive reviews of DRL applied to flow control [59, 61, 40, 20] for more detail. This rapid development has been made possible largely thanks to both the availability of open source DRL frameworks that make it easy to deploy advanced DRL algorithms [25, 18], and open source computational fluid dynamics (CFD) tools [1, 22]. These are being increasingly merged into turn-key frameworks that make it easy to start applying DRL to CFD simulations [65].
@@
In relation to the works described above, several associated challenges have been solved over the last few years. For example, Rabault and Kuhnle [37] have leveraged a multi-environment DRL framework where several numerical simulations run in parallel, each providing its own set of $S_{t},\ A_{t}$ and $R_{t}$ to a common agent, therefore scaling up the number of experiences. With their approach, they showed that speedups in training are obtained proportional to the number of simultaneous environments used. The study by Belus et al. [8] introduced a multi-agent RL (MARL) framework for the control of flow systems with distributed inputs and outputs. MARL was since then repeatedly proved to be effective in problems where the dynamics of the system are translationally invariant in space and multiple actuators are required to control the system, i.e., the control dimensionality is large. With a single-agent RL (SARL) framework, this large dimensionality would result in the curse of dimensionality on the action dimension making learning impractical. In the MARL framework, each agent is defined by separate streams of states, actions and rewards, each unique to a single agent. Moreover, critically, all agents share a common parameterisation, i.e., the same neural network weights. Each agent observes a small subdomain of the full domain, which we define as a ‘pseudo-environment’ (Sec. 2.3), and is thus responsible for only a subset of the total number of control actions, thereby overcoming the curse of dimensionality. The use of a common agent parameterisation is justified by the fact that the governing equations (and thus the qualitative features of the flow dynamics) are invariant over all pseudo-environments, and experience can be shared among MARL agents. Following Belus et al. [8], recent works that employ MARL and have demonstrated its effectiveness are Guastoni et al. [19] for the control of 3D turbulent channels, Suárez et al. [45, 54] to reduce drag over a 3D cylinder, Novati et al. [31], Bae and Koumoutsakos [5] for turbulence closure modelling and Vignon et al. [62] for the control of 2D RBC, to name a few.
@@
The first application of DRL to RBC was presented by Beintema et al. [3] in a 2D domain using the single-agent RL framework. They studied the control of RBC in a square domain confined by no-slip walls with the lateral walls being adiabatic, top wall being isothermal, and temperature control modulation imposed on the bottom wall. The $\mathrm{Ra}_{c}$ at which convection occurs was reduced as the DRL controller converted a single-cell configuration in the uncontrolled system to a vertically stacked double cell configuration, reducing the effective $\mathrm{Ra}$. In a following study by Vignon et al. [62], the MARL framework was used on a domain with periodic boundary conditions on the lateral walls, featuring two convection cells in the uncontrolled state. Given that MARL is advantageous for problems with distributed inputs and outputs for which invariants are observed, a superior performance compared to SARL was demonstrated, in terms of faster learning. MARL achieved a $22.7\%$ reduction in $\mathrm{Nu}$, employing a control strategy that coalesces the two convection cells into a single-cell state, thus reducing the convection intensity. By extension to Beintema et al. [3] and Vignon et al. [62], in the present study, the RBC system is three-dimensional. Therefore, the control dimensionality (number of actuators) and state space dimensionality is even larger than in 2D, and the dynamics are more complex than in 2D. This further motivates the need for the MARL framework for the execution of control in 3D. As far as the authors’ knowledge of the past literature goes, the present study is the first application of MARL to 3D RBC.
@@
The rest of the paper is organised as follows. Sec. 2 provides a detailed description and formulation of the RBC problem (Sec. 2.1), the numerical methods used for the CFD (Sec. 2.2), and a description of the DRL control (Sec. 2.3) and proportional control (Sec. 2.4) methods. In the results section, Sec. 3, the uncontrolled states are described in Sec. 3.1, and results from the training are discussed in 3.2. The evaluation of the trained agent in deterministic mode is provided in Sec. 3.3 and results of deterministic evaluation in a larger domain are provided in Sec. 3.5. Proportional control is discussed in Sec. 3.4. We conclude the paper in Sec. 4. Multimedia (video) files are referred to in the caption of various figures in the text and information about these is provided in the ‘Supplementary Information’ section in Appendix A. We release the full code used in this study as open source materials, and the relevant information is provided in Appendix B.
@@ 2.2 Numerical Method for the CFD
The code is implemented using the open-source spectral Galerkin framework ‘shenfun’ [29], where the equations can be automatically discretized through a high-level scripting language closely resembling the Mathematics. The Navier–Stokes solver has been verified by reproducing the growth of the most unstable eigenmode of the Orr–Sommerfeld equations over long time integrations. The Navier–Stokes and Rayleigh–Bénard solvers are distributed as part of the shenfun software, and there is a demonstration guide published in the documentation [30], which also provides a much more detailed description of the numerical method. Links are provided in Appendix B, and we refer the reader interested in all the technicalities of the numerical solver implementation to the resources detailed therein.
@@ 2.3 DRL Control Methodology
In the following, we use a MARL setup as illustrated in Fig. 2. In MARL, $\mathrm{N}_{\mathrm{ag}}$ multiple agents are each assigned to different local sections of the entire domain, which we refer to as ‘pseudo-environments’. We define a pseudo-environment as the block of the domain directly above each individual segment. Each agent is associated with its own stream of state, action and reward from its corresponding pseudo-environment which we denote as $s^{(j)}_{t}$, $a^{(j)}_{t}$ and $r^{(j)}_{t}$ respectively, $j=1,\cdots,\mathrm{N}_{\mathrm{ag}}$ being the agent index. Here, note the modification in notation from uppercase letters to lowercase letters: uppercase representing states, actions and rewards corresponding to the entire domain as per the SARL framework, and lowercase representing states, actions and rewards specific to each pseudo-environment as per the MARL framework. We represent segment-specific values with the superscript $j$ in parenthesis. Note that the $r^{(j)}_{t}$ is a ‘weighted’ local reward, which is a combination of a ‘local’ and ‘global’ $\mathrm{Nu}$ as described below. This is based on the fact that the RBC environment is invariant in the horizontal directions $x$ and $z$ and so maximising a local reward can correspond to maximising the global reward in a cumulative manner. The state, action and reward tuples from each pseudo-environment and then from each timestep are batched together consecutively up to a size corresponding to the batch size, before an agent update is performed.
@@
Figure 2: An illustration of the MARL setup. The environment consists of the RBC system (top) with a projection of the bottom wall boundary (below) also shown. The bottom boundary is divided into $\mathrm{N}_{\mathrm{ag}}$ square segments where temperature control actuations are applied. Observation probes are distributed uniformly throughout the domain, but for illustration purposes are depicted with black dots in blocks of the domain directly above four pseudo-environments. Each pseudo-environment along with their streams of actions, states and rewards form multiple agents. The index $j$ represents the $j^{\mathrm{th}}$ agent, with $j=1,\dots,\mathrm{N}_{\mathrm{ag}}$. All agents share the same neural network parameterisation in the PPO algorithm.
@@
The states, rewards and actions described above are implemented within the framework of the ‘Tensorforce’ library [26]. Tensorforce contains several optimisation algorithms for reinforcement learning. In the present work, the widely used policy gradient method, proximal policy optimisation (PPO), is employed to train the agent. PPO is an actor-critic method where the actor approximates the policy distribution and the critic approximates the value function of the current state [47]. Both the policy and the critic are represented by neural networks. We make a distinction in terminology between two types of episodes used in the MARL framework: ‘CFD episodes’ and ‘MARL episodes’. A CFD episode is a simulation of the entire RBC domain for 200 actuations. Each actuation in a CFD episode comprises $\mathrm{N_{ag}}$ actions applied locally by each MARL pseudo-environment. Hence, each pseudo-environment applies 200 actuations in one CFD episode. We define 200 actuations applied by a single pseudo-environment as one MARL episode. For $\mathrm{N_{ag}}$ pseudo-environments, one CFD episode equals $\mathrm{N_{ag}}$ MARL episodes. All results in this paper are plotted with respect to CFD episodes only. Therefore, hereafter in the text and figures, the usage of the term ‘episode’ refers to ‘CFD episode’, unless otherwise specified. A set of other training hyperparameters used in this study is provided in Table 1.
In the present work, 200 and 400 total episodes are simulated for $\mathrm{Ra}=500$ and $\mathrm{Ra}=750$ respectively, and each actuation in any given episode is applied for a duration of 15 timesteps. In order to ensure that the agent is also trained on states from longer simulations, we ensure that $20\%$ of the episodes start from the end of the previous episode, and the remaining $80\%$ starts from the end state of the baseline. In this way, the system also explores trajectories spanning a longer time across successive episodes durations, and thus the agent observes more states, including states representative of the long-term behavior of the system. We use a batch size of 3 CFD episodes (i.e., $1923\times\mathrm{N}_{\mathrm{ag}}=3\times 64=$ 192 MARL episodes). We found that with lower batch sizes, training is largely unstable. We explain this observation as follows. Neural networks learn best from data that is uncorrelated. Since data from adjacent MARL environments is strongly (spatially) correlated, a batch size large enough must be chosen such that independent trajectories can be sampled during training. On the other hand, a batch size too large would make training sluggish. Other agent hyperparameters and CFD parameters are provided in Table 1. Training is performed with a discount factor of $\gamma=0.99$. For details of the algorithmic implementation of MARL, we refer the interested reader to Vignon et al. [62]. We follow the same code implementation therein.

### v2

@@ 1 Introduction
Reinforcement learning (RL) is a data-driven approach to the problem of goal-directed sequential decision making. The RL framework involves an ‘agent’, which is its decision-making component, that learns a mapping from states to actions, called the ‘policy’, via a series of trial-and-error interactions with the environment. A reward signal, which is a scalar measure of performance of the environment in relation to the pre-defined goal, is fed back to the agent and is used in the optimization. The state $S_{t}$, action $A_{t}$ and reward $R_{t}$ form the three fundamental channels of communication between the agent and the environment. Each state is assigned a value $V(s)$ which is the expected total reward that can be obtained by that state in the long run, i.e., $V(s)=\operatorname*{\mathbb{E}}\left[\Sigma_{k=0}^{\infty}\gamma^{k}R_{t+k+1}|s\right]$ [48]. The goal of the RL problem, then, is to maximise expected cumulative rewards or ‘returns’. In Deep RL (DRL), function approximators such as neural networks are used to parameterise either the policy, the value function, or both, which are updated during training via interaction with the environment in order to maximise the returns. Various solution methods have been developed to perform this optimisation, namely: (i) value-based methods such as Q-learning method [59, 29], where the value function is approximated and actions are chosen based on the approximated value function, (ii) policy-based methods [51], such as the policy gradient method, where the optimal policy is learned directly, and (iii) actor-critic methods such as the proximal policy optimisation (PPO) method [54], where the critic network computes an advantage function to determine how good an action is compared to the average of all possible actions, given a state. Then, the actor network is updated based on the value of this advantage function. During training, interaction between the agent and the environment occurs on an episode basis, where an ‘episode’ is a single trajectory of the environment for a given fixed duration. Training is performed through trial-and-error that involves a stochastic exploration of the environment’s states, i.e., selecting actions other than those dictated by the latest updated policy, which can improve the estimation of the state values. After training, the agent is evaluated on the environment in a deterministic manner according to the optimal policy, without exploration.
@@
As a consequence, DRL for AFC has been the focus of much work over the last few years. One of the first uses of DRL in AFC was for the stabilisation of the vortex street in the flow past a two-dimensional (2D) cylinder at a low Reynolds number ($Re$) of 100 [39, 38]. Here, the Reynolds number is given as $Re=\tilde{U}\tilde{H}/\tilde{\nu}$, where the symbols $\tilde{U}$, $\tilde{H}$ and $\tilde{\nu}$ are the mean flow velocity, characteristic length scale of the system and kinematic viscosity respectively, and tilde $\tilde{\cdot}$ represents dimensional quantities. With two synthetic jets as the actuators, the drag was reduced by $8\%$ and the learned control strategy was found to reduce pressure drop across the cylinder by increasing the area of the separated region. Following this study, DRL was applied to different configurations of the 2D cylinder. In Tang et al. [57], four synthetic jets were used as DRL controllers for a wider $Re$ range, $Re=100-400$, and a maximum drag reduction of $\approx 38\%$ was achieved. Drag reduction was observed for unseen values of $Re$ as well, showing the robustness of the DRL control process. The work by Ren et al. [40] extends these studies to nonlinear unsteady conditions at $Re=1000$, showing a reduction in drag by $30\%$ along with a reduction of turbulent fluctuations. Varela et al. [64] shows that a transition in the actuation strategy is obtained when the Reynolds number is further increased and turbulent stresses play a dominating role. Fan et al. [17] and Xu et al. [69] use a pair of smaller downstream rotating cylinders as actuators with a combined objective of drag reduction over the main cylinder and also to reduction of power loss due to cylinder rotation. More recently, an application to three-dimensional (3D) cylinders was studied in the $Re$ range of $100-400$ [46], where again, DRL successfully showed a reduction of drag of $\approx 10\%$ for $Re=400$. A number of studies have also considered questions related to the optimization of the full DRL control setup, including probe placement [35, 34, 70] and the use of time history from the state probes [67]. Moreover, applications related to vortex induced vibration [42, 12, 45], or hydrodynamic stealth [43], have also attracted interest. Aside from the cylinder, other applications include the control of chaotic flows such as the flow arising from the Kuramoto–Sivashinsky (KS) equation [9, 68], drag reduction in 3D channel flows [20, 52], suppression of the instability of a one-dimensional (1D) falling liquid film [8], suppression of 2D Rayleigh–Bénard convection (RBC) [3, 63] and other practical problems [42, 43, 28]. We refer the interested reader to some comprehensive reviews of DRL applied to flow control [60, 62, 41, 21] for more detail. This rapid development has been made possible largely thanks to both the availability of open source DRL frameworks that make it easy to deploy advanced DRL algorithms [26, 19], and open source computational fluid dynamics (CFD) tools [1, 23]. These are being increasingly merged into turn-key frameworks that make it easy to start applying DRL to CFD simulations [66].
@@
In relation to the works described above, several associated challenges have been solved over the last few years. For example, Rabault and Kuhnle [38] have leveraged a multi-environment DRL framework where several numerical simulations run in parallel, each providing its own set of $S_{t},\ A_{t}$ and $R_{t}$ to a common agent, therefore scaling up the number of experiences. With their approach, they showed that speedups in training are obtained proportional to the number of simultaneous environments used. The study by Belus et al. [8] introduced a multi-agent RL (MARL) framework for the control of flow systems with distributed inputs and outputs. MARL was since then repeatedly proved to be effective in problems where the dynamics of the system are translationally invariant in space and multiple actuators are required to control the system, i.e., the control dimensionality is large. With a single-agent RL (SARL) framework, this large dimensionality would result in the curse of dimensionality on the action dimension making learning impractical. In the MARL framework, each agent is defined by separate streams of states, actions and rewards, each unique to a single agent. Moreover, critically, all agents share a common parameterisation, i.e., the same neural network weights. Note that here, we consider one of several possible configurations of MARL. An alternative MARL configuration, is where each agent has different individual parameterisations, and they are used to learn policies for different sub-tasks which contribute to a common goal. We refer the interested reader to the details of several such MARL variants in Albrecht et al. [15]. In the present work, each agent observes a small subdomain of the full domain, which we define as a ‘pseudo-environment’ (Sec. 2.3), and is thus responsible for only a subset of the total number of control actions, thereby overcoming the curse of dimensionality. The use of a common agent parameterisation is justified by the fact that the governing equations (and thus the qualitative features of the flow dynamics) are invariant over all pseudo-environments, and experience can be shared among MARL agents. Following Belus et al. [8], recent works that employ MARL and have demonstrated its effectiveness are Guastoni et al. [20] for the control of 3D turbulent channels, Suárez et al. [46, 55] to reduce drag over a 3D cylinder, Novati et al. [32], Bae and Koumoutsakos [5] for turbulence closure modelling and Vignon et al. [63] for the control of 2D RBC, to name a few.
@@
The first application of DRL to RBC was presented by Beintema et al. [3] in a 2D domain using the single-agent RL framework. They studied the control of RBC in a square domain confined by no-slip walls with the lateral walls being adiabatic, top wall being isothermal, and temperature control modulation imposed on the bottom wall. The $\mathrm{Ra}_{c}$ at which convection occurs was reduced as the DRL controller converted a single-cell configuration in the uncontrolled system to a vertically stacked double cell configuration, reducing the effective $\mathrm{Ra}$. In a following study by Vignon et al. [63], the MARL framework was used on a domain with periodic boundary conditions on the lateral walls, featuring two convection cells in the uncontrolled state. Given that MARL is advantageous for problems with distributed inputs and outputs for which invariants are observed, a superior performance compared to SARL was demonstrated, in terms of faster learning. MARL achieved a $22.7\%$ reduction in $\mathrm{Nu}$, employing a control strategy that coalesces the two convection cells into a single-cell state, thus reducing the convection intensity. By extension to Beintema et al. [3] and Vignon et al. [63], in the present study, the RBC system is three-dimensional. Therefore, the actuators are now two-dimensional surfaces at the bottom of the 3D fluid domain, as opposed to one-dimensional segments at the bottom of a 2D domain. As a consequence, in order to match the necessity to cover a 2D surface with actuators, the number of actuators is increased from 10 in Vignon et al. [63] to $8\times 8=64$ in the present case. If we consider a single-agent RL approach, this would mean that the action produced by the DRL agent would be a vector of 64 values, leading to the curse of dimensionality [8, 63]. This further motivates the need for the MARL framework for the execution of control in 3D. As far as the authors’ knowledge of the past literature goes, the present study is the first application of MARL to 3D RBC.
@@
The rest of the paper is organised as follows. Sec. 2 provides a detailed description and formulation of the RBC problem (Sec. 2.1), the numerical methods used for the CFD (Sec. 2.2), and a description of the DRL control (Sec. 2.3) and proportional control (Sec. 2.4) methods. In the results section, Sec. 3, the uncontrolled states are described in Sec. 3.1, and results from the training are discussed in 3.2. The evaluation of the trained agent in deterministic mode is provided in Sec. 3.3 and results of deterministic evaluation in a larger domain are provided in Sec. 3.5. Proportional control is discussed in Sec. 3.4. We conclude the paper in Sec. 4. Multimedia (video) files are referred to in the caption of various figures in the text and information about these is provided in the ‘Supplementary Information’ section in Appendix B. We release the full code used in this study as open source materials, and the relevant information is provided in Appendix C.
@@ 2.2 Numerical Method for the CFD
The code is implemented using the open-source spectral Galerkin framework ‘shenfun’ [30], where the equations can be automatically discretized through a high-level scripting language closely resembling the Mathematics. The Navier–Stokes solver has been verified by reproducing the growth of the most unstable eigenmode of the Orr–Sommerfeld equations over long time integrations. The Navier–Stokes and Rayleigh–Bénard solvers are distributed as part of the shenfun software, and there is a demonstration guide published in the documentation [31], which also provides a much more detailed description of the numerical method. Links are provided in Appendix C, and we refer the reader interested in all the technicalities of the numerical solver implementation to the resources detailed therein.
@@ 2.3 DRL Control Methodology
In the following, we use a MARL setup as illustrated in Fig. 2. In MARL, $\mathrm{N}_{\mathrm{ag}}$ multiple agents are each assigned to different local sections of the entire domain, which we refer to as ‘pseudo-environments’. We define a pseudo-environment as the block of the domain directly above each individual segment. Each agent is associated with its own stream of state, action and reward from its corresponding pseudo-environment which we denote as $s^{(j)}_{t}$, $a^{(j)}_{t}$ and $r^{(j)}_{t}$ respectively, $j=1,\cdots,\mathrm{N}_{\mathrm{ag}}$ being the agent index. Here, note the modification in notation from uppercase letters to lowercase letters: uppercase representing states, actions and rewards corresponding to the entire domain as per the SARL framework, and lowercase representing states, actions and rewards specific to each pseudo-environment as per the MARL framework. We represent segment-specific values with the superscript $j$ in parenthesis. Note that the $r^{(j)}_{t}$ is a ‘weighted’ local reward, which is a combination of a ‘local’ and ‘global’ $\mathrm{Nu}$ as described below. This is based on the fact that the RBC environment is invariant in the horizontal directions $x$ and $z$, so maximising a local reward can correspond to maximising the global reward in a cumulative manner. The state, action and reward tuples from each pseudo-environment and then from each timestep are batched together consecutively up to a size corresponding to the batch size, before an agent update is performed.
@@
Figure 2: An illustration of the MARL setup. The environment consists of the RBC system (top) with a projection of the bottom wall boundary (below) also shown. The bottom boundary is divided into $\mathrm{N}_{\mathrm{ag}}$ square segments where temperature control actuations are applied. Observation probes are distributed uniformly throughout the domain, but for illustration purposes are depicted with black dots in blocks of the domain directly above four pseudo-environments. Each pseudo-environment along with their streams of actions, states and rewards form multiple agents. The index $j$ represents the $j^{\mathrm{th}}$ agent, with $j=1,\dots,\mathrm{N}_{\mathrm{ag}}$. All agents share the same neural network parameterisation in the PPO algorithm. A schematic of two networks are shown representing the actor and critic parameterisation.
@@
The states, rewards and actions described above are implemented within the framework of the ‘Tensorforce’ library [27]. Tensorforce contains several optimisation algorithms for reinforcement learning. In the present work, the widely used policy gradient method, proximal policy optimisation (PPO), is employed to train the agent. PPO is an actor-critic method where the actor approximates the policy distribution and the critic approximates the value function of the current state [48]. Both the policy and the critic are represented by neural networks.
We define an episode as a simulation spanning 200 actuations, each actuation lasting for 15 time steps. At this point, we make a distinction in terminology between two types of episodes used in the MARL framework: ‘CFD episodes’ and ‘MARL episodes’. Both are simulations spanning 200 actuations, however, the difference lies in the part of the domain considered in the simulation. A CFD episode is a simulation of the entire domain involving 200 sets of $\mathrm{N_{ag}}$ actuations, one set for each action step. A MARL episode considers only the 200 actuations as applied locally by a single pseudo-environment. Thus, the simulation of a single CFD episode implies that $\mathrm{N_{ag}}$ MARL episodes are simulated. All results in this paper are plotted with respect to CFD episodes only. Therefore, hereafter in the text and figures, the usage of the term ‘episode’ refers to ‘CFD episode’, unless otherwise specified. A set of other training hyperparameters used in this study is provided in Table 1.
In the present work, 200 and 400 total CFD episodes are simulated for $\mathrm{Ra}=500$ and $\mathrm{Ra}=750$ respectively, and each actuation in any given episode is applied for a duration of 15 timesteps. In order to ensure that the agent is also trained on states from longer simulations, we ensure that $20\%$ of the episodes start from the end of the previous episode, and the remaining $80\%$ starts from the end state of the baseline. In this way, the system also explores trajectories spanning a longer time across successive episodes durations, and thus the agent observes more states, including states representative of the long-term behavior of the system. We use a batch size of 3 CFD episodes (i.e., $1923\times\mathrm{N}_{\mathrm{ag}}=3\times 64=$ 192 MARL episodes). We found that with lower batch sizes, training is largely unstable. We explain this observation as follows. Neural networks learn best from data that is uncorrelated. Since data from adjacent MARL environments is strongly (spatially) correlated, a batch size large enough must be chosen such that independent trajectories can be sampled during training. On the other hand, a batch size too large would make training sluggish. Other agent hyperparameters and CFD parameters are provided in Table 1. Training is performed with a discount factor of $\gamma=0.99$. For details of the algorithmic implementation of MARL, we refer the interested reader to Vignon et al. [63]. We follow the same code implementation therein.